A screw jack has a two-start thread of pitch 5mm. An effort of 40N is applied tangentially to the tommy bar, at a radius of 350 mm, to lift a load of 2200 N.


  1. Calculate the velocity ratio and the mechanical advantage of the jack.
  2. Calculate the work done in overcoming friction when the load is raised a distance of 75 mm.
  3. Evaluate the efficiency of the screw jack and discuss what effect it has on the likelihood of overhauling and the law of the machine.
1
Expert's answer
2021-04-29T07:27:30-0400

Pitch, "p=5\space mm=5\times10^{-3}\space m"

Radius, "r=350\space mm=350\times10^{-3}\space m"

Effort, "E=40\space N"

Load, "L=2200\space N"


(1) Velocity Ratio, "(V.R)=\dfrac{2\pi r}{p}=\dfrac{2\times3.14\times350\times10^{-3}}{5\times10^{-3}}=439.6"

Mechanical Advantage, "(MA)=\dfrac{Load}{Effort}=\dfrac{2200}{40}=55"


(2) Height raised, "h=75\space mm=75\times10^{-3}\space m"

Efficiency, "\eta=\dfrac{MA}{VR}=\dfrac{55}{439.6}\times100=12.511\%"

Work Done "=\dfrac{mgh}{\eta}=\dfrac{2200\times75\times10^{-3}}{0.125}=1320\space J"


(3) "Efficiency, \space \eta=12.511\%"